What is the greatest integer solution of |2x+3|-4 less than or equal to 0?
Answer by stanbon(75887)

What is the greatest integer solution of |2x+3|-4 less than or equal to 0?
------
|2x+3| <= 4
-----
-4 <= 2x+3 <= 4
-----
-7 <= 2x <=1
----
-7/2 <= x <= 1/2
---
The greatest integer solution would be zero.
==================
Cheers,
Stan H.
================